The ends π‘₯ = 0 and π‘₯ = 10 of a thin aluminum bar (𝛼 2 = 0.86) are kept at 0 π‘œπΆ, while the surface of the bar is insulated. Find an expression for the temperature 𝑒(π‘₯,𝑑) in the bar if initially


𝑒(π‘₯, 0) = { 10π‘₯ 0 < π‘₯ < 5 10(10 βˆ’ 10π‘₯) 5 ≀ π‘₯ < 10Β 


1
Expert's answer
2021-10-29T03:01:42-0400

Solution;

x=0 ; x=10

The heat equation is;

"\frac{\delta u}{\delta t}=\alpha^2\frac{\delta^2u}{\delta x^2}......(1)"

The general solution of the heat equation is in the form;

"u(x,t)=(Acos\lambda x+Bsin\lambda x)e^{-\delta ^2\lambda^2t}.......(2)"

The boundary conditions are;

"(i)u(0,t)=0" ; "t\geq0"

"(ii)u(10,t)=0" ; "t>0"

"(iii)u(x,0)=f(x)=10x" ; "0<x<5" ;"10(10-10x);5\leq x<10"

Applying condition (i) in (2),we have;

"0=Ae^{-0.86\lambda^2t}"

Hence;

"A=0"

Equation (2) reduces to ;

"u(x,t)=Bsin\lambda xe^{-0.86\lambda^2t}...(3)"

Apply condition (ii) in (3);

"0=Bsin10\lambda e^{-0.86\lambda^2t}"

"i.e" "10\lambda=nπ" ;"\lambda=\frac{nπ}{10}" (n is an integer)

Substituting in (3);

"u(x,t)=Bsin\frac{nπx}{10}e^{-0.86\frac{n^2π^2}{100}t}"

The most general solution is;

"u(x,t)=\displaystyle\sum_{n=1}^{\infin}B_nsin\frac{nπx}{10}e^{\frac{-0.86n^2π^2}{100}t}" .......(4)

Apply condition (iii) in (4);

"u(x,0)=\displaystyle\sum_{n=1}^{\infin}B_nsin\frac{nπx}{10}=f(x)"

The LHS series is the half range Fourier sine series of the RHS functions.

Therefore;

"B_n=\frac{2}{10}[\int_0^5f(x)sin\frac{nπx}{10}dx+\int_5^{10}f(x)sin\frac{nπx}{10}dx]"

"B_n=\frac{2}{10}[\int_0^510xsin\frac{nπx}{10}dx+\int_5^{10}10(10-10x)sin\frac{nπx}{10}dx]"

Integrate and simplify to obtain;

"B_n=\frac{(90x-100)cos\frac{nπx}{10}}{πn}-\frac{900sin\frac{nπx}{10}}{n^2π^2}"

Replace Bn in (3) ;

"u(x,t)=(\frac{(90x-100)cos\frac{nπx}{10}}{πn}-\frac{900sin\frac{nπx}{10}}{n^2π^2})e^{\frac{-0.86n^2π^2t}{100}}"